# BoostStep: Boosting mathematical capability of Large Language Models via improved single-step reasoning

### v2

@@ Abstract
Cutting-edge large language models (LLMs) demonstrate promising performance in solving complex math problems with a divide-and-conquer pipeline and the assistance of in-context learning (ICL) examples. However, their potential for improvement is limited by two critical problems within their ICL examples: granularity-mismatch and the ensuing negative-effect noise problem. Specifically, the LLMs are capable of the dividing process yet mostly fail by inaccurate reasoning within a few conquer steps, while the ICL examples retrieved in question-grained sometimes lack relevant steps for a specific challenging reasoning step. Further, this disconnect may hinder the correct reasoning due to its irrelevance. To this end, we focus on improving the reasoning quality within each step and present BoostStep. BoostStep aligns the granularity between the retrieving and reasoning on step-grained and provides highly related ICL examples for each reasoning step with a novel ‘first-try’ strategy. BoostStep provides more relevant examples than the coarse question-grained strategy, enhancing the model reasoning quality within each step steadily. BoostStep is a general and robust reasoning-enhancing method that not only improves standalone reasoning performance but also integrates seamlessly with Step-level tree search methods to refine both candidate generation and decision-making. Quantitatively, it improves GPT-4o and Qwen2.5-Math-72B by 3.6% and 2.0% respectively on various mathematical benchmarks, and 7.5% gain combined with tree search.
@@
footnotetext: * indicates equal contribution $\dagger$  indicates corresponding author
@@ 1 Introduction
Figure 1: Our step-level in-context learning outperforms traditional problem-level few-shot learning for about 4% across in-domain, out-domain and cross-modality mathematical benchmark. Moreover, on benchmarks with lower similarity with the example problem set (i.e. OlympiadBench and multi-modal benchmarks), where problem-level few-shot learning may have a negative impact, our strategy still provides valuable guidance.
@@
When solving complex mathematical problems, two capabilities are necessary, which are divide and conquer. The former decomposes the complex problem into several simpler steps to simplify the task, while the latter guarantees the correctness of the one-step reasoning.
@@
Through the analysis of error cases, we found that the current SOTA models have been relatively clear in the divide part, that is, the model can know exactly what tasks should be completed in each step. However, there will still be a lot of mistakes in the conquer section, such as wrong formula use, wrong calculation, insufficient enumeration, etc. To quantitatively substantiate this conclusion, we provide GPT-4o-mini with a ground truth reasoning process to determine whether the error in another response was due to an overarching flawed reasoning approach or a deviation within a particular step. In less advanced models like LLaMA-3.1-8B-Instruct Dubey et al. (2024), 8.7% of errors originate from an overall flawed reasoning approach. In more advanced models like GPT-4o, only 0.8% of errors are ascribable to the overall approach, with the majority resulting from inaccuracies at individual steps of the reasoning process. Therefore, the correctness of one-step reasoning is the bottleneck of reasoning capability.
@@
Various approaches have been employed to improve reasoning correctness, such as producing chains of thought through prompt engineering Kojima et al. (2022); Wei et al. (2022), fine-tuning with mathematical data Shao et al. (2024); Yang et al. (2024); Ying et al. (2024), or generating multiple candidate reasoning paths using Tree Search Methods Zhang et al. (2024b, a).
@@
Among those techniques, in-context learning is a particularly important one, which offers similar examples to provide detailed guidance. However, the examples retrieved by traditional problem-level in-context learning are listed before the reasoning process ever begins, thereby lacking fine-grained guidance during the reasoning process. Moreover, since the example problem can’t be identical to the new one, the irrelevant steps in those examples may even become a distraction from the current reasoning, thus even negatively affecting the conquer capability for some specific steps.
@@
To this end, we refine in-context learning from problem-level to step-level granularity to offer similar example steps during an ongoing reasoning process for fine-grained step-level guidance, as well as ensuring that the introduced example is still highly relevant at the step level to avoid distractions.
@@
Firstly, we have constructed an example problem bank with step-level granularity based on reasoning content, instead of commonly adopted grammatical separation. This ensures the steps in the question bank are consistent with the actual reasoning steps, thereby providing more appropriate guidance.
@@
Building on the step-level granularity within the example problem bank, we propose an approach that incorporates in-context learning through a "first-try" format during an ongoing reasoning process. Specifically, for a given problem to be solved, we break down the solving process into step-by-step reasoning paths. During the reasoning of a single step, we first allow the model to attempt a ‘first try’ to comprehend what the model currently needs to reason about. Based on this initial attempt, we searched the question bank to find similar steps that can guide the model to accurately output the current step. This first-try strategy helps ensure a higher similarity between the retrieved examples and the current step. Therefore, the distraction from irrelevant steps can be avoided and the guidance effect is improved.
@@
Compared with traditional in-context learning in problem-level granularity, our method is capable of providing examples during the reasoning process directly based on the steps to be solved, thereby offering more relevant guidance. Our approach demonstrates significant improvements over traditional few-shot learning across various open-source benchmarks, with an average increase of 3.6%.
@@
Moreover, our method also reduces the dependency on the similarity between the example dataset and the target problem since two entirely different problems can still share similar key steps, such as utilizing prime factorization for both calculating the least common multiple and determining divisibility. Consequently, dissimilar problems can still offer effective guidance. For example, on multi-modal benchmarks with lower similarity to the example problem bank like MathVision Wang et al. (2024), traditional few-shot learning can even have negative effects. In contrast, our approach continues to achieve significant improvements.
@@
Our method significantly enhances single-step reasoning accuracy through step-level in-context learning, making it highly compatible with various current reasoning strategies that employ step-level tree search. Typically, a tree-search method requires a reasoning model to generate multiple step-wise candidate reasoning paths and a critic model to evaluate the correctness of these candidates. Our approach can be integrated into both aspects in a plug-and-play manner. Specifically, when the reasoning model generates new candidate reasoning nodes, our method can introduce similar examples in the aforementioned ‘first-try’ manner to improve the accuracy of candidates. Additionally, it can aid the critic model by incorporating similar correct reasoning steps into the evaluation of candidate reasoning processes to provide similar guidance. Experimental results indicate that both applications contribute positively to the model’s reasoning capabilities. Detailed illustration will be shown in Sec. 3.4
@@ 2 Related Works
Stepwise Mathematical Reasoning. Recently, to further enhance mathematical reasoning capabilities, many studies have shifted the granularity of mathematical reasoning from the problem level to the step level. This approach involves addressing each next step individually and completing small segments of reasoning within the overall task. These works often employ tree searching strategies like Tree of Thoughts (ToT) Yao et al. (2024) or Monte Carlo Tree Search Zhang et al. (2024b, a); Chen et al. (2024); Feng et al. (2023); Zhu et al. (2022), extending multiple steps to optimize step answers and ultimately obtain the optimal solution. Additionally, Process Supervision Models (PRMs) Lightman et al. (2023); Luo et al. (2024) are frequently used to verify the correctness of new candidate nodes in real-time and prune reasoning paths, thereby improving the accuracy of the final answer. This more detailed auxiliary strategy demonstrates greater potential.
In-context Learning in Mathematical Reasoning.  In-context learning can provide low-cost guidance to models through similar examples, thereby enhancing the quality of model outputs and their ability to follow prompts. Consequently, it has been widely adopted. However, research on in-context learning within mathematical reasoning tasks remains insufficient. Typically, this approach involves providing the model with similar problems and their ground truth solutions to offer a general strategy for solving new problems Hendrycks et al. (2021); Wei et al. (2022). Some efforts have been made to improve the relevance of retrieved examples by designing better retrieval mechanisms and incorporating appropriate reference rejection techniques Liu et al. (2024b). Others try to provide high-level context instead to improve the generalizability Wu et al. (2024). However, all these methods share a common limitation: the lack of fine-grained real-time guidance at the step level.
@@ 3.1 Revisiting In-Context Learning from Conditional Probability
To this end, we propose step-level in-context learning and a first-try strategy to provide detailed and relevant example steps when in step-level reasoning. Specifically, when generating new steps $s_{i+1}$ based on previous reasoning steps $s_{i},s_{i-1},\ldots,s_{1}$ and question $q$, we first utilize a first-try strategy to obtain an approximate estimate of $s^{first}_{i+1}$. Then, we use this $s^{first}_{i+1}$ to retrieve a similar step $s^{\prime}_{n+1}$ along with the corresponding $q^{\prime},s^{\prime}_{1},s^{\prime}_{2},\ldots,s^{\prime}_{n}$. Since these two steps are similar, a very reasonable assumption is that $P(s^{\prime}_{n+1}\mid q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n})$ closely approximates $P(s_{{gt}_{i+1}}\mid q,s_{1},\ldots,s_{i})$. Therefore, the generated step $s_{i+1}=\arg\max\limits_{s}P_{model}(s\mid q,s_{1},\ldots,s_{i},q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n},,s^{\prime}_{n+1},)$ will be more closed to $s_{{gt}_{i+1}}$ comparing to $s^{0-shot}_{i+1}$. Details about our step-level in-context learning and first-try strategy will be explained in Sec. 3.3
@@ 3.2 Step-Level Example Problem Bank
Some approaches Lightman et al. (2023) proposed using the period ’.’ as a clear semantic delimiter to segment steps for training PRM. This strategy allows for the quick decomposition of each step from a complete process without any additional assistance. However, this simple decomposition mode is obviously unreliable. Essentially, a single reasoning step should have a consistent target and encompass a complete thought process, making it the atomic granularity of reasoning. Using a period ’.’ as a delimiter may disrupt this atomicity. For example, it may split a complete enumeration for the same objective into multiple steps.
Therefore, we suggest that the most appropriate method for step segmentation is to allow the reasoning model itself to autonomously decompose the process. This approach ensures that the granularity of the decomposed steps in example problem bank aligns with that of the real-time reasoning steps. Specifically, we define the concept of a step through prompts, which encapsulate a complete and simple inference. This guides GPT-4o in decomposing the answer at the step level.
@@
* The proposed step-level example problem bank is available at https://github.com/beichenzbc/BoostStep
@@ 3.3 Step-Level ICL with First-try Strategy
At the more granular step level, however, the situation becomes much more complex. A simple strategy is to perform retrieval using the given problem and all preceding reasoning steps $s_{i-1},s_{i-2},\ldots,s_{1},q$. The clear drawback of this method is the excessive length of the retrieval content, which diminishes the emphasis on the uniqueness of the current step. Another strategy is to use the previous step $s_{i-1}$ to retrieve $s^{\prime}_{j-1}$ from a step-level database, thereby guiding the reasoning of $s_{i}$ through the correct resolution of $s^{\prime}_{j}$. However, this approach is rather crude, as it models stepwise reasoning as a Markov process, which is evidently unreasonable. Similar steps can be applicable to different reasoning tasks, and therefore similarity in the previous step does not necessarily indicate that the retrieved subsequent step will provide valuable guidance for the reasoning in the current step.
@@ 3.4 Step-Level Guidance in Tree Search
Generally, tree search methods necessitate two key components: a reasoning model that generates step-level reasoning and a Process-Supervision Reward Model (PRM) that continuously evaluates the current reasoning step in real time. Our method is beneficial for both of these components. It enhances the step-level reasoning performed by the reasoning model and improves the effectiveness of the PRM in evaluating current reasoning steps.
For the reasoning model, tree search methods inherently require step-by-step reasoning expansion. When expanding at node $s_{i}$, we simply apply the previously mentioned strategy: the model performs $n$ first tries to obtain $n$ example steps. For each example, the model then completes the reasoning to generate $n$ child nodes $s^{1}_{i+1},s^{2}_{i+1},\ldots,s^{n}_{i+1}$ with the help of these examples. Our strategy improves the accuracy of individual child nodes $s^{j}_{i+1}$, thereby enhancing the overall reasoning quality.
Evidently, judgment ability is closely related to reasoning ability. Therefore, since our strategy can enhance the accuracy of single-step reasoning, a reasonable assumption is that introducing appropriate example steps can similarly improve the PRM’s ability to assess the correctness of the current reasoning process. In particular, when evaluating the correctness of an inference step candidate $s^{j}_{i}$, we retrieve similar steps $s^{\prime}_{k}$ along with their corresponding preceding steps $s^{\prime}_{k-1},\ldots,s^{\prime}_{1}$ and question $q^{\prime}$ from the step-level example bank. Similarly, the probability distributions $P(s^{\prime}_{k}|s^{\prime}_{k-1},\ldots,s^{\prime}_{1},q^{\prime})$ and $P(s_{gt_{i}}|s_{i-1},\ldots,s_{1},q)$ exhibit similarities. This resemblance aids in assessing the discrepancy between $s^{j}_{i}$ and $s_{gt_{i}}$, thereby enhancing the accuracy of the critic model’s evaluations.
Detailed ablation experiments in Sec. 4.5 demonstrate that both strategies contribute positively to the overall reasoning quality of step-level tree search methods.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated impressive ability in solving complex mathematical problems with multi-step reasoning and can be further enhanced with well-designed in-context learning (ICL) examples. However, this potential is often constrained by two major challenges in ICL: granularity mismatch and irrelevant information. We observe that while LLMs excel at decomposing mathematical problems, they often struggle with reasoning errors in fine-grained steps. Moreover, ICL examples retrieved at the question level may omit critical steps or even mislead the model with irrelevant details. To address this issue, we propose BoostStep, a method that enhances reasoning accuracy through step-aligned ICL, a novel mechanism that carefully aligns retrieved reference steps with the corresponding reasoning steps. Additionally, BoostStep incorporates an effective "first-try" strategy to deliver exemplars highly relevant to the current state of reasoning. BoostStep is a flexible and powerful method that integrates seamlessly with chain-of-thought (CoT) and tree search algorithms, refining both candidate selection and decision-making. Empirical results show that BoostStep improves GPT-4o’s CoT performance by 4.6% across mathematical benchmarks, significantly surpassing traditional few-shot learning’s 1.2%. Moreover, it can achieve an additional 7.5% gain combined with tree search. Surprisingly, it enhances state-of-the-art LLMs to solve challenging math problems using simpler examples. It improves DeepSeek-R1-671B’s performance on AIME by 2.2%, leveraging simple examples only from the MATH dataset.
@@
00footnotetext: * indicates equal contribution $\dagger$  indicates corresponding author
@@ 1 Introduction
Figure 1: Our step-aligned in-context learning (ICL) outperforms traditional problem-level few-shot learning for about 4% across in-domain, out-domain and cross-modality mathematical benchmark on GPT4o. Moreover, on benchmarks with lower similarity with the reference problem set (i.e. OlympiadBench and multi-modal benchmarks), where problem-level ICL may have a negative impact, BoostStep still provides valuable guidance.
@@
When solving complex mathematical problems, cutting-edge LLMs often adopt a multi-step reasoning strategy. Specifically, they first decompose a complex problem into several simpler steps and then solve each single step independently.
@@
Through the analysis of error cases, we found that current SOTA models are relatively correct in the step-dividing phase, that is, the model can know exactly what tasks should be completed in each step. However, there are still a lot of mistakes within each reasoning step, such as wrong formula use, wrong calculation, insufficient enumeration, etc. To quantitatively substantiate this observation, we provide GPT-4o-mini with a ground truth reasoning process to determine whether the error in another response was due to an overarching flawed reasoning approach or a deviation within a particular step. In less advanced models like LLaMA-3.1-8B Dubey et al. (2024), 91.3% of errors originate from single-step reasoning. In more advanced models like GPT-4o, up to 99.2% of errors are ascribable to some particular steps. This exaggerated proportion suggests that the correctness of single-step reasoning is the bottleneck of reasoning capability.
@@
Various approaches have been employed to improve reasoning correctness, such as producing chains of thought through prompt engineering Kojima et al. (2022); Wei et al. (2022), fine-tuning with mathematical data Shao et al. (2024); Yang et al. (2024); Ying et al. (2024), or generating multiple candidate reasoning paths using Tree Search Methods Zhang et al. (2024b, a); Wang et al. (2024b).
@@
Among those techniques, in-context learning is a particularly important one, which offers similar examples to provide detailed guidance. However, the examples retrieved by traditional problem-level in-context learning are listed before the reasoning process, thereby lacking fine-grained guidance during the reasoning process. Moreover, since the example problem can’t be identical to the new one, the irrelevant steps in those examples may even become a distraction from the current reasoning, thus even negatively affecting the single-step reasoning capability for some specific steps.
@@
To this end, we refine in-context learning from problem-level to step-level granularity to offer similar example steps during an ongoing reasoning process for fine-grained step-aligned guidance. We also ensure that the introduced example is still relevant at the step level to avoid distractions.
@@
Firstly, we have constructed an example problem bank with step-level granularity based on reasoning content instead of commonly adopted grammatical separation. This ensures the steps in the problem bank are consistent with the actual reasoning steps, thereby providing more appropriate guidance.
@@
Building on the step-level granularity within the example problem bank, we propose an approach that incorporates in-context learning through a "first-try" format during an ongoing reasoning process. Specifically, for a given problem to be solved, we break down the solving process into step-by-step reasoning paths. During the reasoning of a single step, we first allow the model to attempt a ‘first try’ to comprehend what the model currently needs to reason about. Based on this initial attempt, we searched the problem bank to find similar steps that can guide the model to accurately output the current step. This helps ensure a higher similarity between the retrieved examples and the current step so the distraction from irrelevant steps can be avoided and the guidance effect can be improved.
@@
Compared with traditional problem-level ICL, our method provides examples during the reasoning process directly based on the steps to be solved, thereby offering more relevant guidance. It demonstrates significant improvements over traditional few-shot learning across various benchmarks, with an average increase of 3.4% on GPT-4o.
@@
Moreover, our method also reduces the sensitivity to the similarity between the example and the target problem, as two different problems can still share similar steps. Consequently, dissimilar problems can still offer effective guidance. On multi-modal benchmarks with lower similarity to example problems, traditional few-host learning has a detrimental effect, resulting in an accuracy reduction of 0.9% on GPT-4o. In contrast, our approach still achieves an improvement of 2.8%.
Besides, BoostStep also shows a promising potential to improve the reasoning quality on harder problems with simpler examples. With examples from MATH Hendrycks et al. (2021), it helps Deepseek-R1 achieve an improvement of 2.2% on the much more challenging AIME problems.
@@
Moreover, our method is also highly compatible with various current reasoning strategies that employ step-level tree search. Typically, a tree-search method requires a reason model to generate multiple step-level candidate reasoning paths and a critic model to evaluate the correctness of these candidates. Our approach can be integrated into both aspects. Specifically, when the reason model generates new candidate reasoning nodes, our method can introduce similar examples in the aforementioned ‘first-try’ manner to improve the accuracy of candidates. Additionally, it can aid the critic model by incorporating similar example steps into the evaluation of candidate reasoning processes to provide similar guidance. Experiments indicate that both applications contribute positively and bring about an improvement of 8.5% jointly on GPT-4o.
@@ 2 Related Works
Step-level Mathematical Reasoning. Recently, to further enhance mathematical reasoning capabilities, many studies have shifted the granularity of mathematical reasoning from the problem level to the step level. This approach involves addressing each next step individually and completing small segments of reasoning within the overall task. These works often employ tree searching strategies like Tree of Thoughts (ToT) Yao et al. (2024); Besta et al. (2024) or Monte Carlo Tree Search Zhang et al. (2024b, a); Chen et al. (2024); Feng et al. (2023); Zhu et al. (2022), extending multiple steps to optimize step answers and ultimately obtain the optimal solution. Additionally, Process Supervision Models (PRMs) Lightman et al. (2023); Luo et al. (2024) are frequently used to verify the correctness of new candidate nodes in real-time and prune reasoning paths, thereby improving the accuracy of the final answer. This more detailed auxiliary strategy demonstrates greater potential.
In-context Learning in Mathematical Reasoning.  In-context learning can provide low-cost guidance to models through similar examples, thereby enhancing the quality of model outputs and their ability to follow prompts. Consequently, it has been widely adopted. However, research on in-context learning within mathematical reasoning tasks remains insufficient. Typically, this approach involves providing the model with similar problems and their ground truth solutions to offer a general strategy for solving new problems Hendrycks et al. (2021); Wei et al. (2022). Some efforts have been made to improve the relevance of retrieved examples by designing better retrieval mechanisms and incorporating appropriate reference rejection techniques Liu et al. (2024b). Others try to provide high-level context instead to improve the generalizability Wu et al. (2024). However, all these methods share a common limitation: the lack of fine-grained step-level guidance. Some recent approaches Dong et al. (2024) introduce ICL into the reason process. However, they still perform ICL in problem granularity and thus may not offer effective guidance for next-step reasoning.
@@ 3.1 Revisiting In-Context Learning from Conditional Probability
To this end, we propose step-aligned in-context learning and a first-try strategy to provide detailed and relevant example steps when in step-level reasoning. Specifically, when generating new steps $s_{i+1}$ based on previous reasoning steps $s_{i},s_{i-1},\ldots,s_{1}$ and question $q$, we first utilize a first-try strategy to obtain an approximate estimate of $s^{first}_{i+1}$. Then, we use this $s^{first}_{i+1}$ to retrieve a similar step $s^{\prime}_{n+1}$ along with the corresponding $q^{\prime},s^{\prime}_{1},s^{\prime}_{2},\ldots,s^{\prime}_{n}$. Since these two steps are similar, a very reasonable assumption is that $P(s^{\prime}_{n+1}\mid q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n})$ closely approximates $P(s_{{gt}_{i+1}}\mid q,s_{1},\ldots,s_{i})$. Therefore, the generated step $s_{i+1}=\arg\max\limits_{s}P_{model}(s\mid q,s_{1},\ldots,s_{i},q^{\prime},s^{\prime}_{1},\ldots,s^{\prime}_{n},,s^{\prime}_{n+1},)$ will be more closed to $s_{{gt}_{i+1}}$ comparing to $s^{0-shot}_{i+1}$. Details about our step-level in-context learning and first-try strategy will be explained in Sec. 3.3
@@ 3.2 Step-Level Example Problem Bank
Some approaches Lightman et al. (2023) proposed using a clear semantic delimiter like the period ’.’ or a new line to segment steps. This strategy allows for the quick decomposition of each step from a complete process without any additional assistance. However, this simple decomposition mode is obviously unreliable. Essentially, a single reasoning step should have a consistent target and a complete thought process, making it the atomic granularity of reasoning. Using a period ’.’ or a new line as a delimiter may disrupt this atomicity. For example, it may split a complete enumeration for the same objective into multiple steps.
Therefore, we suggest that the most appropriate method for step segmentation is to allow the reason model itself to autonomously decompose the process. This approach ensures that the granularity of the decomposed steps in example problem bank aligns with that of the real-time reasoning steps. Specifically, we define the concept of a step through prompts, which encapsulate a complete and simple inference. This guides GPT-4o in decomposing the answer at the step level.
@@
111* The proposed step-level example problem bank is available at https://github.com/beichenzbc/BoostStep
@@ 3.3 Step-Level ICL with First-try Strategy
At the more granular step level, however, the situation becomes much more complex. A simple strategy is to perform retrieval using the given problem and all preceding reasoning steps $s_{i-1},s_{i-2},\ldots,s_{1},q$. The clear drawback of this method is the excessive length of the retrieval content, which diminishes the emphasis on the uniqueness of the current step. Another strategy is to use the previous step $s_{i-1}$ to retrieve $s^{\prime}_{j-1}$ from a step-level database, thereby guiding the reasoning of $s_{i}$ through the correct resolution of $s^{\prime}_{j}$. However, this approach is rather crude, as it models step-level reasoning as a Markov process, which is evidently unreasonable. Similar steps can be applicable to different reasoning tasks, and therefore similarity in the previous step does not necessarily indicate that the retrieved subsequent step will provide valuable guidance for the reasoning in the current step.
@@ 3.4 Step-Level Guidance in Tree Search
Generally, tree search methods necessitate two key components: a reason model that generates step-level reasoning and a Process-Supervised Reward Model (PRM) that continuously evaluates the current reasoning step in real time. Our method is beneficial for both of these components. It enhances the step-level reasoning performed by the reason model and improves the effectiveness of the PRM in evaluating current reasoning steps.
For the reason model, tree search methods inherently require step-by-step reasoning expansion. When expanding at node $s_{i}$, we can apply the previously mentioned strategy: the model performs $n$ first tries and retrieve for $n$ example steps. For each example, the model then completes the reasoning to generate $n$ child nodes $s^{1}_{i+1},\ldots,s^{n}_{i+1}$ with the help of these examples. Similarly, our strategy can improve the accuracy of individual nodes $s^{j}_{i+1}$.
Evidently, judgment ability is closely related to reasoning ability. Therefore, since our strategy can enhance the accuracy of single-step reasoning, a reasonable assumption is that introducing appropriate example steps can improve the PRM’s ability to assess the correctness of the current reasoning process. In particular, when evaluating the correctness of an inference step candidate $s^{j}_{i}$, we retrieve similar steps $s^{\prime}_{k}$ along with their corresponding preceding steps $s^{\prime}_{k-1},\ldots,s^{\prime}_{1}$ and question $q^{\prime}$ from the step-level example bank. Similarly, the probability distributions $P(s^{\prime}_{k}|s^{\prime}_{k-1},\ldots,s^{\prime}_{1},q^{\prime})$ and $P(s_{gt_{i}}|s_{i-1},\ldots,s_{1},q)$ exhibit similarities. This resemblance aids in assessing the discrepancy between $s^{j}_{i}$ and $s_{gt_{i}}$, thereby enhancing the accuracy of the critic model’s evaluations.
Detailed ablation experiments in Sec. 4.5 demonstrate that both strategies contribute positively to step-level tree search methods.